OpenStudy (anonymous):

IK how to do this i have done 13 others but I can't seem to solve this one A number is divided by 3. If 8 is subtracted from the quotient you get -6. What is the number? A. -6 B. 3 C. 6 D. 42

OpenStudy (anonymous):

Let the number be x. Then your equation is x/3-8=-6.

OpenStudy (anonymous):

You have to solve that for x.

OpenStudy (anonymous):

So add 8 on both sides that gives you x/3=-6+8=2

OpenStudy (anonymous):

Now multiply by 3 giving you x=2*3=6, so C is the right answer.

OpenStudy (saifoo.khan):

Oh Yeah!